Would you please show the steps in how to compute these? This as you know is z scores. Thank you.
The average commute time via train from the Chicago O'Hare Airport to downtown is 60 minutes with a standard deviation of 15 minutes. Assume that the commute times are normally distributed. What proportion of commutes would be:
a. longer than 80 minutes?
b. less than 50 minutes?

Answer by stanbon(75887)

The average commute time via train from the Chicago O'Hare Airport to downtown is 60 minutes with a standard deviation of 15 minutes. Assume that the commute times are normally distributed. What proportion of commutes would be:
a. longer than 80 minutes?
z(80) = (80-60)/15 = 20/15 = 4/3
P(x > 80) = P(z > 4/3) = normalcdf(4/3,100) = 0.0912
---------------------
b. less than 50 minutes?
z(50) = (50-60)/15 = -10/15 = -2/3
P(x < 50) = P(z < -2/3) = normalcdf(-100,-2/3) = 0.2525
===========================
Cheers,
Stan H.
====================